OpenStudy (anonymous):

Find the center,vertices,co-vertices,foci,and asymptoes for the hyperbola: x^2/4-y^2/9=1

OpenStudy (mertsj):

Did you draw the graph?

OpenStudy (mertsj):

\[\frac{(x-h)^2}{a^2}-\frac{(y-k)^2}{b^2}=1\]

OpenStudy (mertsj):

Center is (h,k)

OpenStudy (anonymous):

i dont have nything to draw the graph wth

OpenStudy (mertsj):

Do you have a graphing calculator? Or a piece of paper and a pencil?

OpenStudy (anonymous):

paper and penical yes

OpenStudy (mertsj):

So make a sketch of it. You can get pretty close.

OpenStudy (mertsj):

What is the center?

OpenStudy (mertsj):

Does it open up and down or right and left?

OpenStudy (anonymous):

(0,0) is the center

OpenStudy (mertsj):

What is a^2